Sir, how to integrate dx/[k-m{(2nx)^1/2}] from 0 to x k,m&n are constants

Sir,


how to integrate


dx/[k-m{(2nx)^1/2}]


from 0 to x


k,m&n are constants

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear aman

 oxdx/[k-m2nx]

we can also wrote

 

I = 1/m2n oxdx/[(k/m2n)-x]

    =  1/m2n oxdx/[a-x]            where a =(k/m2n)

 I= 1/m2n ox[a+x]dx/[a-x]

  I= 1/m2n ox[a]dx/[a-x]  +  1/m2n ox[x]dx/[a-x]

first part u can easly calculate

for second part  

 I2 = oxxdx/x[a-x]

put x = t

         dx = 2tdt

I2 = ox2t2dt/[a-t2]

this is also a standard form

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin